# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v3

@@ 1. Vacuum pump (any model capable of producing 500 mbar vacuum)
An Example of Chemical Protocol Procedure Design System Message: You are an excellent expert in experimental protocol design. Given a user requirement for the experiment, and the materials that may be required, your task is to design the procedure of the experiment. Do not output any other characters. User Message: Below is a user’s experimental design requirement: I am exploring the ability to separate enantiomers of racemic mixtures accurately and efficiently. My purpose is to utilize capillary electrophoresis (CE) with a novel cationic cyclodextrin, mono-6-deoxy-6-(3R,4R-dihydroxypyrrolidine)-$\beta$-CD chloride (dhypy-CDCl), as a chiral selector. I aim to determine if this specific single isomer cyclodextrin can enhance the enantioseparation process in CE, thereby providing a potential improvement over current methods in terms of resolution and selectivity. The goal is to validate this method for effective use in chiral analysis and possibly apply it to pharmaceuticals where the correct enantiomer is crucial for desired therapeutic effects. Give the materials, reagents, and equipment that may be required for the experiment, please design a detailed and step-by-step procedure for the experiment based on the user’s needs: 1. dhypy-CDCl 2. Sodium hydroxide (NaOH) 3. Sodium phosphate monobasic dehydrate (NaH2PO4•2H2O) 4. HPLC-grade methanol (MeOH) 5. HPLC-grade acetonitrile (ACN) 6. Ultra-pure water 7. 2-(3-Chlorophenoxy)propionic acid (3-ClPOP A) 8. Dansyl-DL-valine cyclohexylammonium salt (Dns-Val) 9. Beckman P/ACE MDQ CE system 10. 50 cm x 50 $\mu$m I.D. uncoated fused-silica capillary 11. 32 Karat Software (Version 5.0) Expected Answer: 1. Add a certain amount of NaH2PO4•2H2O into ultra-pure water to prepare 50 mM NaH2PO4 stock solutions. 2. Adjust the pH of NaH2PO4 solutions using 1 M NaOH until the desired pH (generally 6 9) was obtained. 3. Add a certain amount of CD selector into the NaH2PO4 solution to obtain the desired CD concentration (generally 0 30 mM). 4. Add a certain amount of organic modifier (MeOH or ACN) (generally 0 20%, vol/vol) into the solutions from Step 3. 5. Transfer 1.5 mL solutions into two 2 mL-vials. 6. Degas the buffer before use. 7. Add a certain amount of racemic analytes into 50/50 (vol/vol) methanol/water mixture to form stock solutions of 50 $\mu$g•mL-1. 8. Transfer 1.5 mL stock solutions into 2 mL-vials. 9. Degas the analyte solution before use. 10. Flush the capillary using 1 M NaOH solution for 30 min. 11. Flush the capillary using 0.1 NaOH solution for 30 min. 12. Flush the capillary using ultra-pure water for 30 min. 13. Flush the capillary using running buffer for 15 min. 14. Put the cartridge with fused-silica capillary on the CE equipment and put the two buffer vials and one analyte vial in the sample trays. 15. Set sample injection by pressure at 0.5 psi for 4 s. 16. Set the separation voltage as 15 kV. 17. Start injection and separation. 18. Collect separation data and stop running.
Appendix A7 Data Quality
To ensure the generated data quality, we employed a three-stage data screening process:
1. Initial screening by LLMs. Our primary concern is the ”Multiple Choice Questions (MCQ)” tasks entirely generated by LLMs, such as the Literature QA task. To ensure the correctness of LLM-generated answers, we first explicitly instructed LLM during data generation that the correct options must be clearly identifiable from the provided literature snippets. After data generation, we prompt GPT-4o (Table A10) to simulate an open-book exam task, where it determines whether each question’s answer could be found in the corresponding literature snippet, and if so, GPT-4o could provide answers to questions based on text snippets (for example, identifying the correct option for multiple-choice questions from a snippet). By comparing these answers with the previously generated answers, we can verify the accuracy of the original answers. Otherwise, we consider the answers to the questions unverifiable, and we simply delete them.
2. Human evaluation. We randomly selected approximately 5% of the questions from each task and provided them with two experts in biology and chemistry, with the assistance of five graduate students from related fields. It took a week to complete the quality evaluation. During the evaluation, we used the instructions in Table A11 to guide the evaluators, asking them to thoroughly assess the data and classify it into binary categories of ”Yes” and ”No” for quality. Only data that fully met the requirements was rated ”Yes.” Ultimately, we identified 2.1% instances of data that were rated ”No”.
3. Post-screening by LLMs. We employed LLMs to summarize the failure types of these low-quality entries, and added them into the prompt to conduct a full dataset quality assessment, discarding similar types of low-quality questions.
We repeated the stage2&3 twice, and finally the low-quality entries identified by experts is less than 0.2%. By implementing these stages, we ensure that the SciKnowEval dataset maintains a high standard of data quality.
Appendix A8 Data Privacy and Copyright
The development of the SciKnowEval dataset involves rigorous procedures to ensure data privacy protection and compliance with copyright.
Data Privacy. All data is thoroughly screened for any potentially sensitive information. Identifiable data, such as personal names, institutional affiliations, and proprietary research details, are anonymized or removed. This step ensures that the dataset can be used without risking the exposure of private information.
Data Copyright. We enforce strict copyright review procedures for all documents and data used in our dataset. This involves verifying the sources of the data and ensuring that all content is either in the public domain or used with proper authorization. We obtain necessary permissions and licenses for proprietary content to prevent any infringement of intellectual property rights. Additionally, proper attribution is given to all sources of data. We ensure that the dataset includes clear citations and references to original works. Additionally, we respect the licensing terms of third-party data and comply with any restrictions or requirements specified by the content creators.
@@
Below is a piece of text and a multiple-choice question, your task is to determine whether the question stems from this text and whether the correct answer to the question can be found within the text. If the text explicitly mentions the content being asked in the question, and the answer to the question is also in the text, then output ”Yes” followed by a space and the letter of the correct option, e.g., ”Yes A”. Otherwise, output ”No”.
@@
Your output should be ”Yes” followed by a space and the letter of the correct option if the question stems from the text and the correct answer can be found within the text. Otherwise, output ”No” only.
@@
Your output should be ”Yes” if the question and the answer is high-quality and ”No” otherwise.
Full Disclosure
Prior to participation, all experts are provided with a comprehensive explanation of the experiment and their informed consent is obtained. They are explicitly informed that their annotation results will be utilized to refine our benchmark dataset.
Confidentiality and Privacy
To guarantee that no experts suffers any adverse effects from their involvement, all collected data and annotation outcomes are strictly for scientific research purposes. All information is maintained in strict confidence, ensuring that no personal details of the experts are disclosed and that there are no negative repercussions for them.
@@
We designed scoring prompts for LLMs to evaluate some generation tasks, including text summary, reagent & procedure generation, and so on.
@@
Prompt for Evaluating Text Summary System Message: You are an assistant proficient in generating text summaries. Given a text, its summary, and a model-generated summary, your task is to score the model-generated summary based on the content of the text and its summary. The score ranges from 1 to 5, where 1 means the summary content is very poor, and 5 means the generated summary is of equally high quality as the text’s summary in terms of coherence, relevance, information retention, fluency, conciseness, and usefulness. User Message: Below is a text, its summary, and a model-generated summary. Your task is to score the model-generated summary based on the content of the text and its summary. [Text Start]: {question} [Text End] [Text Summary Start]: {answer} [Text Summary End] [Generated Summary Start]: {response} [Generated Summary End] You should strictly follow the following criteria for scoring: 1. Coherence: You need to judge whether the generated summary is logically consistent and assess the fluency of its internal structure. 2. Relevance: You need to strictly judge whether the generated summary is closely related to the content of the text. 3. Information retention: You need to carefully judge whether the summary contains key information from the text, such as crucial terms, characters, relationships, etc. 4. Fluency: You need to determine whether the language of the summary is smooth enough. 5. Conciseness: You need to focus on whether the summary is sufficiently concise and clear, and whether it does not include any unnecessary information. 6. Usefulness: You need to judge whether the generated summary enables readers to quickly understand the original text. You can refer to the provided text’s summary and, by comparing the two summaries, further assess the quality of the generated summary. Your should directly output the ”Rating: Score” only, e.g. ”Rating: 1” for poor and ”Rating: 5” for excellent. Do not output any other characters. Your output is:
@@
Prompt for Evaluating Reagent Generation System Message: You are a scientific assistant proficient in experimental protocol design. Given a question about reagent selection, a correct reagent selection plan, and a model-generated answer, your task is to score the model-generated answer based on the question and the correct reagent selection plan. The score ranges from 1 to 5, where 1 indicates that the answer is very poor, and 5 indicates that the generated answer effectively addresses the question and matches well with the correct reagent selection. User Message: Below is a question about reagent selection, a correct reagent selection plan, and a model-generated answer. Your task is to score the model-generated answer based on the question and the correct reagent selection plan. [Question Start]: {question} [Question End] [Correct Plan Start]: {answer} [Correct Plan End] [Generated Answer Start]: {response} [Generated Answer End] You should strictly follow the following criteria for scoring: 1. Relevance: You need to assess whether the generated answer is closely related to the provided reagent selection question. 2. Logic and Coherence: You need to evaluate whether the generated answer is logically and structurally correct and coherent. 3. Usefulness: You need to carefully judge whether the generated answer can truly solve or partially solve the provided question. A useful answer should propose feasible solutions. 4. Detail: You need to rigorously judge whether the generated answer contains detailed information, including the specific names of experimental reagents and materials, specific dosages, and concentrations used. 5. Correctness: You need to focus on comparing the generated answer with the provided correct reagent selection plan. Only the reagents and materials that appear in the correct reagent selection plan can be considered correct. Please consider these criteria comprehensively when scoring. Your should directly output the ”Rating: Score” only, e.g. ”Rating: 1” for poor and ”Rating: 5” for excellent. Do not output any other characters. Your output is:

### v4

@@ 1. Vacuum pump (any model capable of producing 500 mbar vacuum)
[…]
Appendix A8 Prompts for Evaluating Data Quality
@@
Below is a piece of text and a multiple-choice question, your task is to determine whether the question stems from this text and whether the correct answer to the question can be found within the text. If the text explicitly mentions the content being asked in the question, and the answer to the question is also in the text, then output "Yes" followed by a space and the letter of the correct option, e.g., "Yes A". Otherwise, output "No".
@@
Your output should be "Yes" followed by a space and the letter of the correct option if the question stems from the text and the correct answer can be found within the text. Otherwise, output "No" only.
@@
Your output should be "Yes" if the question and the answer is high-quality and "No" otherwise.
@@
We designed scoring prompts for LLMs to evaluate some generation tasks, including text summary, reagent & procedure generation, and so on. Notably, we incorporated reference answers into each prompt to assist with the evaluation. We emphasize that using a powerful proprietary model to rate responses based on these reference answers makes the evaluation results relatively reliable [16].
@@
Prompt for Evaluating Text Summary System Message: You are an assistant proficient in generating text summaries. Given a text, its summary, and a model-generated summary, your task is to score the model-generated summary based on the content of the text and its summary. The score ranges from 1 to 5, where 1 means the summary content is very poor, and 5 means the generated summary is of equally high quality as the text’s summary in terms of coherence, relevance, information retention, fluency, conciseness, and usefulness.
User Message:
Below is a text, its summary, and a model-generated summary. Your task is to score the model-generated summary based on the content of the text and its summary.
[Text Start]:
{question}
[Text End]
[Text Summary Start]:
{answer}
[Text Summary End]
[Generated Summary Start]:
{response}
[Generated Summary End]
You should strictly follow the following criteria for scoring:
1. Coherence: You need to judge whether the generated summary is logically consistent and assess the fluency of its internal structure.
2. Relevance: You need to strictly judge whether the generated summary is closely related to the content of the text.
3. Information retention: You need to carefully judge whether the summary contains key information from the text, such as crucial terms, characters, relationships, etc.
4. Fluency: You need to determine whether the language of the summary is smooth enough.
5. Conciseness: You need to focus on whether the summary is sufficiently concise and clear, and whether it does not include any unnecessary information.
6. Usefulness: You need to judge whether the generated summary enables readers to quickly understand the original text.
You can refer to the provided text’s summary and, by comparing the two summaries, further assess the quality of the generated summary.
Your should directly output the "Rating: Score" only, e.g. "Rating: 1" for poor and "Rating: 5" for excellent. Do not output any other characters.
Your output is:
@@
Prompt for Evaluating Reagent Generation System Message: You are a scientific assistant proficient in experimental protocol design. Given a question about reagent selection, a correct reagent selection plan, and a model-generated answer, your task is to score the model-generated answer based on the question and the correct reagent selection plan. The score ranges from 1 to 5, where 1 indicates that the answer is very poor, and 5 indicates that the generated answer effectively addresses the question and matches well with the correct reagent selection.
User Message:
Below is a question about reagent selection, a correct reagent selection plan, and a model-generated answer. Your task is to score the model-generated answer based on the question and the correct reagent selection plan.
[Question Start]:
{question}
[Question End]
[Correct Plan Start]:
{answer}
[Correct Plan End]
[Generated Answer Start]:
{response}
[Generated Answer End]
You should strictly follow the following criteria for scoring:
1. Relevance: You need to assess whether the generated answer is closely related to the provided reagent selection question.
2. Logic and Coherence: You need to evaluate whether the generated answer is logically and structurally correct and coherent.
3. Usefulness: You need to carefully judge whether the generated answer can truly solve or partially solve the provided question. A useful answer should propose feasible solutions.
4. Detail: You need to rigorously judge whether the generated answer contains detailed information, including the specific names of experimental reagents and materials, specific dosages, and concentrations used.
5. Correctness: You need to focus on comparing the generated answer with the provided correct reagent selection plan. Only the reagents and materials that appear in the correct reagent selection plan can be considered correct.
Please consider these criteria comprehensively when scoring. Your should directly output the "Rating: Score" only, e.g. "Rating: 1" for poor and "Rating: 5" for excellent. Do not output any other characters.
Your output is: